Solve the following equations.   3/7 + x = 17/7 
Solve the following equations.   3/7 + x = 17/7 

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

3/7+x=17/7
x = 17/7–3/7
x=14/7
x=2

Thanks